1a) Find the real numbers a and b such that

i5(3 - 2i)2 - (a - bi) = a - bi.


b. Let z1 =  √3 + i and z2 = √3 - i. Show that

z161 + z261 = 261√3.


(Remember that: cos(- θ) = cos(θ), sin(θ) = -sin(θ) = -sin(θ), cos(2kπ + θ) = cos(θ) and sin(2kπ + θ) = sin(θ) )

1
Expert's answer
2022-02-23T13:06:06-0500

1a)


"i^5(3 - 2i)^2 - (a - bi) = a - bi."

"i(9-12i-4)=2a-2bi"

"12+5i=2a-2bi"

"a=6, b=-5\/2"

b.


"z_1=\sqrt{3}+i=2(\cos (\pi/6)+i\sin (\pi/6))"

"z_2=\sqrt{3}-i=2(\cos (-\pi/6)+i\sin (-\pi/6))"

"z_1^{61}=(2(\cos (\pi/6)+i\sin (\pi/6)))^{61}"

"=2^{61}(\cos (61\pi/6)+i\sin (61\pi/6))"

"=2^{61}(\cos (\pi/6)+i\sin (\pi/6))"

"z_2^{61}=(2(\cos (-\pi/6)+i\sin (-\pi/6)))^{61}"

"=2^{61}(\cos (-61\pi/6)+i\sin (-61\pi/6))"

"=2^{61}(\cos (-\pi/6)+i\sin (-\pi/6))"

"=2^{61}(\cos (\pi/6)-i\sin (\pi/6))"

"z_1^{61}+z_2^{61}=2^{61}(\cos (\pi/6)+i\sin (\pi/6))"

"+2^{61}(\cos (\pi/6)-i\sin (\pi/6))"

"=2^{61}(2\cos (\pi/6))=2^{61}\sqrt{3}"